0.632 g of ammonium chloride (NH4Cl) are added to a 1 L solution of water at a temperature of 73°C in a tube separated by a semi-permeable membrane. Calculate the osmotic pressure of the solution.
1
Expert's answer
2020-05-19T08:47:25-0400

posm=CMRT;

T=273+73=346 K

CM=m(NH4Cl)/(Mr(NH4Cl)*V)=0.632/(53.5*1)=0.0118 M;

posm=0.0118*8.314*346=33.94 kPa;

Answer: 33.94 kPa